OpenStudy (anonymous):

Medal and Fan Im not very good at algebra Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach? 1.3 ft 2.6 ft 26.0 ft 27.6 ft

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

Using the same H(t) formula from the last one H(t) = -16t² + vt + s, what's v and what's s?

OpenStudy (anonymous):

v is 42 s is 0?

OpenStudy (anonymous):

yep, so we have H(t) = -16t² + 42t so now we just need to find the vertex. For a quadratic function in the form \[y=ax^2+bx+c\] the vertex is \[\left( \frac{ -b }{ 2a },f \left( \frac{ -b }{ 2a } \right) \right)\] We have a = -16, b = 42, and c = 0 So for the time, find \[t=\frac{ -42 }{ 2*-16 }\] Then plug it into our function to get the max height

OpenStudy (anonymous):

ok so it would be h(t)=-16t^2 + 42t and substitute 1.31 for t?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the answer becomes D correct?

OpenStudy (anonymous):

yes